OpenStudy (juan1857):

Write the polar equation r^2-2rsintheta=0 in rectangular form.

OpenStudy (freckles):

recall r^2=x^2+y^2 r sin(theta)=y r cos(theta)=x tan(theta)=y/x hint: you only need two of these for this one

OpenStudy (juan1857):

so the first two?

OpenStudy (freckles):

yep

OpenStudy (juan1857):

ok so how do you apply them, lol

OpenStudy (freckles):

hint you are trying to write your equation in terms of x and y

OpenStudy (freckles):

r^2... equals? r sin(theta).... equals?

OpenStudy (freckles):